# RFUAV: A Benchmark Dataset for Unmanned Aerial Vehicle Detection and Identification

### v1

@@ 2 RF-based Drone Detection System
Figure 1: Simulation diagram of the actual application of RF-Based UAV monitoring system.
@@ 3 Limitations of Existing Datasets
(ii) A two-stages model for drone detection and identification: We present an open-source RF-based drone detection and identification architecture that leverages mainstream models such as ViT [8], SwinTransformer [14], YOLO [20], ResNet [10], and Faster R-CNN [21], EfficientNet [24], and MobileNet [11].
@@ 4.1 Data Collection
The hardware utilized in this setup primarily consists of the following: (i) two Universal Software Radio Peripheral (USRP) X310 units for signal acquisition and transmission, (ii) a computer equipped with an AMD Ryzen 5950X CPU, 32 GB of RAM, and a 2TB Samsung M.2 Solid State Drive (SSD), running Ubuntu 22.04 Linux and GNU Radio v3.9.10, serving as the signal collector, (iii) a HUAWEI SP310 10GbE card acting as the intermediary connecting the USRP X310 units and the signal collector, (iv) a spectrum analyzer with a frequency range of 9 kHz to 13.2 GHz, used to verify the feasibility of the acquired signals, and (v) a laptop running adobe audition (AU) for signal observation. Initially, we establish a direct ethernet connection between the laptop and the signal collector. Two USRP X310 units are interfaced with the signal collector via 10GbE I/O ports. The first USRP X310 (USRPI) captures raw data, which is subsequently stored directly on the signal collector’s SSD. The second USRP X310 (USRPO) interfaces with a spectrum analyzer to assess the quality of the collected data.
@@
Figure 2: Physical connection diagram of the signal acquisition platform
In total, approximately 1.3 TB of raw data are collected from 37 distinct types of drones. These raw signals encompass Frequency-Hopping Spread Spectrum (FHSS) signals transmitted by the remote control to regulate the drone’s behavior. FHSS signals are characterized by a fixed hopping bandwidth (FHSBW), hopping duration (FHSDT), hopping duty cycle (FHSDC), and hopping pattern period (FHSPP). Additionally, some drones equipped with cameras possess the capability to capture and transmit images. The raw signals from these drones also include fixed bandwidth signals used for image transmission. Typically, these image transmission signals manifest near the frequency bands associated with the drone’s FHSS signals.
@@ 4.2.1 Basic Information about Dataset
Figure 3: Basic information of the original data in the dataset
@@
It is noteworthy that the frequency hopping pattern of certain drones’ FHSS signals exhibit variability contingent upon the drone’s behavior. For example, during the communication establishment process, the remote controller of the WFLY WFT09SII generates FHSS signals characterized by shorter FHSDC and irregular hopping patterns. However, once a stable connection is established between the drone and its controller, and subsequent to transmission from the flight controller to the drone, there is a transition in the FHSS signal towards a longer-period signal. At this stage, the frequency hopping pattern resembles that of a triangular waveform. To ensure comprehensive sample coverage, our dataset encompasses classifications based on various actions performed by the drone. For each individual drone, both the pairing process and flight operations are documented, thereby providing an extensive range of signal types across different operational phases.
@@ 4.2.2 Analysis of the Main Characteristics of Drone Signal
In the practical application of RF-based drone detection systems, signals within the same frequency band as drones also include Bluetooth, WiFi, and various other sources of interference. These interfering signals can significantly hinder the model’s capacity to accurately classify different drone types. To assist readers in identifying more effective solutions, we conduct an analysis of the characteristics of drone signals through their time-frequency representations.
Unlike signals from other sources, drone FHSS signals exhibit distinct characteristics, including fixed FHSBW and FHSDT. The FHSDC and the FHSPP vary across different drone models as shown in Fig. 4.
Figure 4: Main feature statistics of drone signals in the dataset
@@ 4.2.3 Floating Time-Duration of Image Transmission Signal in RFUAV
Figure 5: Main distribution of drone image transmission signal duration in the dataset
@@ 4.4 Signal Separation and Parameter Estimation
RFUAV can also be employed in various signal processing analyses. In this section, we present a signal separation and SNR estimation algorithm that utilizes a subset of signals from RFUAV. This methodology capitalizes on the capabilities of RFUAV data to effectively isolate individual drone communication signals from background noise and interference, while accurately estimating the SNR to assess the quality of the received signals.
We first employ a dual sliding window-based signal detection algorithm in the time domain to identify the starting and ending positions of a signal clip. Based on the results of this separation, we employ Welch’s method to estimate the SNR of each signal. The specific implementation process is detailed Appendix. D.
@@ 5.1.1 Dataset
The raw signal and their SNR-adjusted counterparts are transformed into spectrograms with identical frequency resolutions, employing four distinct CMAP methods (for more detailed CMAP methods analysis, see Appendix. A) to produce images of consistent resolution. The spectrograms of the raw signals and the SNR-adjusted signals at 20 dB serve as training data, whereas the spectrograms of the SNR-adjusted signals with SNR levels ranging from -20 dB to 20 dB (in 2 dB increments) are used as validation data.
@@ 5.1.2 Model Pre-selection
Table 1: Model comparison results
@@ 5.1.3 CMAP Comparison
Table 2: CMAP comparison results
@@
Figure 7: CMAP Comparison Results
As shown in figure and table, we observe that the ViT-L-16 model achieves the highest OA of 58.16% on datasets using the ”Hot” CMAP method. In scenarios where the SNR is between -8 dB and 20 dB, the model consistently achieves the highest precision, with ${\text{OA}}_{\text{SNR}\geq 10}$ at 99.70% and ${\text{OA}}_{-8\leq\text{SNR}\leq 8}$ at 55.80%, respectively. Meanwhile, for datasets using the ”HSV” CMAP method, the ViT model achieves the highest ${\text{OA}}_{\text{SNR}\leq-10}$ of 31.84%. Therefore, the ”Hot” CMAP method should be considered the most suitable option for this task.
@@ 5.2.2 Comparison Result
Table 3: Frequency Resolution Comparison
@@
As depicted in the table and figure, the impact of increasing STFTP on model accuracy demonstrates a boundary effect (for a more detailed analysis, refer to Appendix. D). Once STFTP reaches 256, further increments do not result in significant improvements in model accuracy. Instead, the accuracy gradually stabilizes, with larger STFTP values causing fluctuations in accuracy between 0.12% and 0.05%. Therefore, in this experiment, the model with STFTP set to 256, achieves the highest OA of 58.28%, ${\text{OA}}_{\text{SNR}\leq-10}$ of 20.33%, and ${\text{OA}}_{-8\leq\text{SNR}\leq 8}$ of 56.00%.
@@ Appendix 0.A STFTP Transformation Detail
In this section, we demonstrate effect of the total number of STFT points in the discrete Fourier transform (STFTP) on the quality of the spectrogram.
@@
where, $w[n]$ is the window function at discrete time index $n$, $N$ indicates the STFTP. The Hamming window is typically used to reduce the side lobes in the frequency domain and improve frequency resolution.
From Eq. 1 and 2, it can be observed that for the same signal frame, increasing STFTP results in a higher number of signal samples used for the STFT transformation. This leads to a richer frequency representation. To provide more detailed frequency information in the transformed signal, thereby making the frequency expression of the signal clearer and better reflecting the true frequency characteristics of the signal, we present a 3D spectrogram analysis. We select nine window lengths of $2^{n}$ (where $n=3,4,...,11$). The results are shown in Fig. 89, and 10.
As illustrated in the figure, a smaller STFTP results in a loss of frequency details in the spectrogram. When STFTP is 8, the frequency resolution of the spectrogram is limited to 12.5 MHz, meaning that frequency variations within each 12.5 MHz bandwidth cannot be captured and are represented by a single point. At this setting, only the presence of the image transmission signal at the center frequency is visible, but its boundary is unclear, and the FHSS signal cannot be distinguished. As the STFTP increases to 64, the frequency resolution improves to 1.5625 MHz, allowing the boundary of the 20 MHz bandwidth image transmission signal to be identified, although the FHSS signal remains blurred. When the STFTP reaches 512, the frequency resolution increases to 195.3125 kHz, enabling clear differentiation between the image transmission and FHSS signals. For the spectrogram of the flight control signal from the WFLY ET16S in Fig. 9, when the STFTP is below 64 (i.e., the frequency resolution is below 1.5625 MHz), the broader spectrum components are indistinguishable, and the FHSS signal spectrum is barely discernible. However, once the STFTP increases to 256 and the frequency resolution reaches 390.625 kHz, the spectrum of the FHSS becomes much clearer.
Combining the experiment results presented in Sec. 5.2.2 with the Fig. 7, we can conclude that excessively high frequency resolution is unnecessary and may excessively compromise time resolution. As shown in Fig. 8, when the STFTP reaches 65536, the temporal variations in the FHSS signals begin to blend together. At an STFTP value of 524288, the image transmission signals merge entirely in time domain, becoming indistinguishable.
@@ Appendix 0.B Data Collection Detail
During the collection process, the transmitters of 37 different UAV types are either directly connected to the receiver of the USRPI or positioned close proximity to it, ensuring optimal signal quality of the collected data. The open-source software GNURadio is utilized to configure the acquisition parameters of the USRPI. According to previous research, the communication bandwidth used by civilian drones does not exceed 100 MHz. Therefore, a uniform sampling rate of 100 MHz is set during the data acquisition process. The center frequency of the acquisition is dynamically adjusted based on the specific signals of different drones, ensuring complete capture of the drone signals. Additionally, the gain settings are dynamically adjusted according to the SNR of the original signals to maintain a high SNR for all acquired data. Some DJI drones exhibit variability in their video transmission signal bandwidth (VTS BW), alternating among several predefined bandwidth options. To ensure comprehensive data coverage, multiple acquisition attempts are conducted to capture all potential scenarios, guaranteeing that each specific bandwidth configuration is represented in the dataset.
The primary challenge encounters during data collection is the uncertainty regarding the frequency bands of certain drones. To overcome this issue, we employ spectrogram to manually search for signals across a wide frequency range. Capturing such signals incurs extremely manual costs; however, these efforts have successfully brought these raw drone signals to public attention for the first time, thereby greatly enhancing the value and relevance of the dataset.
@@ Appendix 0.C Dataset Details
Table 4: Summary of basic data of drone signals in the dataset
@@ Appendix 0.D Signal Separation and Parameter Estimation Detail
In this section, we present a signal separation and SNR estimation method using some signals processing techniques. we use an 80 ms segment of the DJI FPV COMBO radio frequency signal clip as sample to showcase the whole process (Fig. 11).
@@
Firstly, we employ the STFT to analyze its distribution in both time and frequency domains. The window function employed in this analysis is the Hamming window, which has a length of 512. the signal sample comprises: (i) a downlink video transmission signal modulated by Orthogonal Frequency Division Multiplexing (OFDM) , with a bandwidth of 40 MHz; (ii) an uplink control signal using Frequency-Hopping Spread Spectrum (FHSS) technology, characterized by a bandwidth of 3 MHz for each hop; and (iii) a data frame containing DroneID information as per the DJI OcuSync protocol [22], with a bandwidth of 10 MHz. Although the frequency bands employed by these three types of signals overlap, they do not coincide temporally. Consequently, we can implement a signal detection algorithm based on dual sliding windows in the time domain to identify both the starting and ending positions of the signal frames. Given that the frame durations for these three types of signals differ significantly, we can classify them based on this distinguishing characteristic. Finally, we perform STFT separately to each classification result to obtain the each components
@@
Based on the separation results, we take the first occurring video transmission signal (denote as $X_{1}(n)$), and set a sliding window with a width of 40MHz on the PSD. Starting from the minimum frequency and moving to the max frequency, and calculate the energy value within the window. When the energy within the window is at its maximum, we determine that the current signal matches the window function, and the frequency corresponding to the center of the window function on the frequency axis is then considered as the frequency of the signal.
Based on the starting and ending positions of the signal, we can calculate the energy of the signal and the noise in the time domain, and then estimate the SNR. Assuming the starting index of $X_{1}(n)$ is ${N_{s}}$ and the ending index is ${N_{e}}$, the length of the $X_{1}(n)$ is $L$. Then, the sum of the signal and noise power can be expressed as

### v2

@@ 2 RF-based Drone Detection System
Figure 1: Schematic representation of the RF-based UAV monitoring framework. The system captures and analyzes radio frequency signals emitted by UAV operators to detect and classify drone activity. The monitoring station processes the received signals, facilitating UAV detection and identification of drone types.
@@ 3 Limitations of Existing Datasets
(ii) A two-stage model for drone detection and identification: We present an open-source RF-based drone detection and identification architecture that leverages mainstream models such as ViT [8], SwinTransformer [14], YOLO [20], ResNet [10], and Faster R-CNN [21], EfficientNet [24], and MobileNet [11].
@@ 4.1 Data Collection
The hardware utilized in this setup primarily consists of the following: (i) two Universal Software Radio Peripheral (USRP) X310 units for signal acquisition and transmission, (ii) a computer equipped with an AMD Ryzen 5950X CPU, 32 GB of RAM, and a 2TB Samsung M.2 Solid State Drive (SSD), running Ubuntu 22.04 Linux and GNU Radio v3.9.10, serving as the signal collector, (iii) a HUAWEI SP310 10GbE card acting as the intermediary connecting the USRP X310 units and the signal collector, (iv) a spectrum analyzer with a frequency range of 9 kHz to 13.2 GHz, used to verify the feasibility of the acquired signals, and (v) a laptop running adobe audition (AU) for signal observation. Initially, two USRP X310 units are interfaced with the signal collector via 10GbE I/O ports. The first USRP X310 (USRPI) captures raw data, which is subsequently stored directly on the signal collector’s SSD. The second USRP X310 (USRPO) interfaces with a spectrum analyzer to assess the quality of the collected data.
@@
Figure 2: Hardware architecture and interconnections of the RF signal acquisition platform. The system integrates multiple components, including a GNU Radio controller, USRP devices, a spectrum analyzer, and an AU monitoring unit. These components facilitate high-quality RF signal collection, processing, and analysis.
In total, approximately 1.3 TB of raw data are collected from 37 distinct types of drones. These raw signals encompass Frequency-Hopping Spread Spectrum (FHSS) signals transmitted by the remote control to regulate the drone’s behavior. FHSS signals are characterized by a fixed hopping bandwidth (FHSBW), hopping duration (FHSDT), hopping duty cycle (FHSDC), and hopping pattern period (FHSPP). Additionally, some drones equipped with cameras possess the capability to capture and transmit images. The raw signals from these drones also include fixed bandwidth signals used for image transmission. Typically, these image transmission signals manifest near the frequency bands associated with the FHSS signals.
@@ 4.2.1 Basic Information about Dataset
Figure 3: Statistical overview of the original dataset. The figure presents key attributes of the collected RF data, including file size (GB), signal-to-noise ratio (SNR) in dB, and center frequency (MHz × $10^{-1}$) for various UAV controllers.
@@
It is noteworthy that the frequency hopping pattern of certain drones’ FHSS signals exhibits variability contingent upon the drone’s behavior. For example, during the communication establishment process, the remote controller of the WFLY WFT09SII generates FHSS signals characterized by shorter FHSDC and irregular hopping patterns. However, once a stable connection is established between the drone and its controller, and subsequent to transmission from the flight controller to the drone, there is a transition in the FHSS signal towards a longer-period signal. At this stage, the frequency hopping pattern resembles that of a triangular waveform. To ensure comprehensive sample coverage, our dataset encompasses classifications based on various actions performed by the drone. For each individual drone, both the pairing process and flight operations are documented, thereby providing an extensive range of signal types across different operational phases.
@@ 4.2.2 Analysis of the Main Characteristics of Drone Signal
In the practical application of RF-based drone detection systems, signals within the same frequency band as drones also include Bluetooth, WiFi, and various other sources of interference. These interfering signals can significantly hinder the model’s capacity to accurately classify different drone types. To assist readers in identifying more effective solutions, we conduct an analysis of the characteristics of drone signals through their time-frequency representations. Unlike signals from other sources, drone FHSS signals exhibit distinct characteristics, including fixed FHSBW and FHSDT. The FHSDC and the FHSPP vary across different drone models as shown in Fig. 4.
Figure 4: Comprehensive statistical analysis of key features observed in drone communication signals. Left Polar Plot: Displays the frequency-hopping signal bandwidth (FHSBW, MHz), image transmission signal bandwidth (VTSBW, MHz), and frequency-hopping signal duration time (FHSDT, ms). Right Polar Plot: Depicts the frequency-hopping duty cycle (FHSDC, ms) and frequency-hopping signal periodicity (FHSPP, ms).
@@ 4.2.3 Floating Time-Duration of Image Transmission Signal in RFUAV
Figure 5: Main distribution of image transmission signal duration in the RFUAV
@@ 4.4 Signal Separation and Parameter Estimation
RFUAV can also be employed in various signal processing analyses. In this section, we present a signal separation and SNR estimation algorithm that utilizes a subset of signals from RFUAV.
We first employ a dual sliding window-based signal detection algorithm in the time domain to identify the starting and ending positions of a signal clip. Based on the results of this separation, we use Welch’s method to estimate the PSD of the signal and determine the center frequency of the signal. Finally, the SNR is calculated in the time domain according to the definition of SNR. The specific implementation process is detailed Appendix. D.
@@ 5.1.1 Dataset
The raw signal and their SNR-adjusted counterparts are transformed into spectrograms with identical frequency resolutions, employing four distinct CMAP methods (for more detailed CMAP methods analysis, see Appendix. A) to produce images of consistent resolution. The spectrograms of the raw signals and the SNR-adjusted signals at 20 dB serve as training data, whereas the spectrograms of the SNR adjusted signals with SNR levels ranging from -20 dB to 20 dB (in 2 dB increments) are used as validation data.
@@ 5.1.2 Model Pre-selection
Table 1: Detailed model comparison results
@@ 5.1.3 CMAP Comparison
Table 2: Detailed CMAP comparison results
@@
Figure 7: (a) presents the comparative results of CMAP. (b) presents the STFTP comparison results,
As shown in the figure and table, we observe that the ViT-L-16 model achieves the highest OA of 58.16% on datasets using the ”Hot” CMAP method. In scenarios where the SNR is between -8 dB and 20 dB, the model consistently achieves the highest precision, with ${\text{OA}}_{\text{SNR}\geq 10}$ at 99.70% and ${\text{OA}}_{-8\leq\text{SNR}\leq 8}$ at 55.80%, respectively. Meanwhile, for datasets using the ”HSV” CMAP method, the ViT model achieves the highest ${\text{OA}}_{\text{SNR}\leq-10}$ of 31.84%. Therefore, the ”Hot” CMAP method should be considered the most suitable option for this task.
@@ 5.2.2 Comparison Result
Table 3: Detailed STFTP comparison results
@@
As depicted in the table and figure, the impact of increasing STFTP on model accuracy demonstrates a boundary effect (for a more detailed analysis, refer to Appendix. D). Once STFTP reaches 256, further increments do not result in significant improvements in model accuracy. Instead, the accuracy gradually stabilizes, with larger STFTP values causing fluctuations in accuracy between 0.05% and 0.12%. Therefore, in this experiment, the model with STFTP set to 256, achieves the highest OA of 58.28%, ${\text{OA}}_{\text{SNR}\leq-10}$ of 20.33%, and ${\text{OA}}_{-8\leq\text{SNR}\leq 8}$ of 56.00%.
@@ Appendix 0.A STFTP Transformation Detail
In this section, we demonstrate the effect of the total number of STFT points in the discrete Fourier transform (STFTP) on the quality of the spectrogram.
@@
where $w[n]$ is the window function at discrete time index $n$, $N$ indicates the STFTP.
From Eq. 1 and Eq. 2, it can be observed that increasing STFTP results in a higher number of signal samples used for the STFT transformation. This leads to a richer frequency representation. To provide more detailed frequency information in the transformed signal, thereby making the frequency expression of the signal clearer and better reflecting the true frequency characteristics of the signal, we present a 3D spectrogram analysis. We select nine window lengths of $2^{n}$ (where $n=3,4,...,11$). The results are shown in Fig. 89, and 10.
As illustrated in the figure, a smaller STFTP results in a loss of frequency details in the spectrogram. When STFTP is 8, the frequency resolution of the spectrogram is limited to 12.5 MHz, meaning that frequency variations within each 12.5 MHz bandwidth cannot be captured and are represented by a single point. In this setting, only the presence of the image transmission signal at the center frequency is visible, but its boundary is unclear, and the FHSS signal cannot be distinguished. As the STFTP increases to 64, the frequency resolution improves to 1.5625 MHz, allowing the boundary of the 20 MHz bandwidth image transmission signal to be identified, although the FHSS signal remains blurred. When the STFTP reaches 512, the frequency resolution increases to 195.3125 kHz, enabling clear differentiation between the image transmission and FHSS signals. For the spectrogram of the flight control signal from the WFLY ET16S in Fig. 9, when the STFTP is below 64 (i.e., the frequency resolution is below 1.5625 MHz), the broader spectrum components are indistinguishable, and the FHSS signal spectrum is barely discernible. However, once the STFTP increases to 256 and the frequency resolution reaches 390.625 kHz, the spectrum of the FHSS becomes much clearer.
Combining the experiment results presented in Sec. 5.2 with Fig. 7 (b), we can conclude that excessively high-frequency resolution is unnecessary and may excessively compromise time resolution. As shown in Fig. 8, when the STFTP reaches 65536, the temporal variations in the FHSS signals begin to blend together. At an STFTP value of 524288, the image transmission signals merge entirely in the time domain, becoming indistinguishable.
@@ Appendix 0.B Data Collection Detail
During the collection process, the transmitters of 37 different UAV types are either directly connected to the receiver of the USRPI or positioned close proximity to it, ensuring optimal signal quality of the collected data. The open-source software GNURadio is utilized to configure the acquisition parameters of the USRPI. According to previous research, the communication bandwidth used by civilian drones does not exceed 100 MHz. Therefore, a uniform sampling rate of 100 MHz is set during the data acquisition process. The center frequency of the acquisition is dynamically adjusted based on the specific signals of different drones, ensuring complete capture of the drone signals. Additionally, the gain settings are dynamically adjusted according to the SNR of the original signals to maintain a high SNR for all acquired data. Some DJI drones exhibit variability in their video transmission signal bandwidth, alternating among several predefined bandwidth options. To ensure comprehensive data coverage, multiple acquisition attempts are conducted to capture all potential scenarios, guaranteeing that each specific bandwidth configuration is represented in the dataset.
The primary challenge encountered during data collection is the uncertainty regarding the frequency bands of certain drones. To overcome this issue, we employ a spectrogram to manually search for signals across a wide frequency range. Capturing such signals incurs extremely manual costs; however, these efforts have successfully brought these raw drone signals to public attention for the first time, thereby greatly enhancing the value and relevance of the dataset.
@@ Appendix 0.C Dataset Details
Table 4: Summary of fundamental characteristics of drone signals in the dataset. The figure presents key statistical attributes, including file size, signal-to-noise ratio (SNR), and center frequency, frequency-hopping signal bandwidth (FHSBW, MHz), image transmission signal bandwidth (VTSBW, MHz), frequency-hopping signal duration time (FHSDT, ms), frequency-hopping duty cycle (FHSDC, ms) and frequency-hopping signal periodicity (FHSPP, ms). providing a comprehensive overview of the collected RF signal data for UAV identification and analysis.”
@@ Appendix 0.D Signal Separation and Parameter Estimation Detail
In this section, we present a signal separation and SNR estimation method using some signal processing techniques. we use an 80 ms segment of the DJI FPV COMBO radio frequency signal clip as a sample to showcase the whole process (Fig. 11).
@@
Firstly, we employ the STFT to analyze its distribution in both time and frequency domains. The window function employed in this analysis is the Hamming window, which set a length of 512. the signal sample comprises: (i) a downlink video transmission signal modulated by orthogonal frequency division multiplexing (OFDM), with a bandwidth of 40 MHz; (ii) an uplink control signal using frequency-hopping spread spectrum (FHSS) technology, characterized by a bandwidth of 3 MHz for each hop; and (iii) a data frame containing DroneID information as per the DJI OcuSync protocol, with a bandwidth of 10 MHz. Although the frequency bands employed by these three types of signals overlap, they do not coincide temporally. Consequently, we can implement a signal detection algorithm based on dual sliding windows in the time domain to identify both the starting and ending positions of the signal frames. Given that the frame durations for these three types of signals differ significantly, we can classify them based on this distinguishing characteristic. Finally, we perform STFT separately for each classification result to obtain each component.
@@
Based on the separation results and prior knowledge, we take the first occurring video transmission signal (denoted as $X_{1}(n)$), and set a sliding window with a width of 40MHz on its PSD. Starting from the minimum frequency to the maximum frequency, and calculating the energy value within the window. When the energy within the window is at its maximum, we determine that the current signal matches the window function, and the frequency corresponding to the center of the window function on the frequency axis is then considered as the frequency of the signal.
Based on the starting and ending positions of the signal, we can calculate the energy of the signal and the noise in the time domain, and then estimate the SNR. Assuming the starting index of $X_{1}(n)$ is ${N_{s}}$ and the ending index is ${N_{e}}$, the length of the $X_{1}(n)$ is $L$. Then, the sum of the signal and noise power can be expressed as: